OpenStudy (anonymous):

If you have Sec 80 what does Csc equal?

OpenStudy (solomonzelman):

\[Sec80=1~/~Cos~(80)\]\[\frac{1}{\sqrt{1-\sin^2x}}\]\[\frac{1}{\sqrt{1-\frac{1}{Sin(80)^2}}}\] I was like what?

OpenStudy (solomonzelman):

Opps hold on.

OpenStudy (solomonzelman):

In the last step it is csc not sin.

OpenStudy (solomonzelman):

\[\frac{1}{\sqrt{1-\frac{1}{Csc^2(80)}}}\]

OpenStudy (solomonzelman):

But I am really not sure if this is good.